An electronic field with a strength of 1.25×10⁴N/C exerted 2.50 Nnon a certain test charges, q. determine charge on q.

1
Expert's answer
2022-05-19T16:40:21-0400
"F=qE"

"q=\frac{F}{E}=\frac{2.50}{1.25*10^{4}}=2.0*10^{-4}\:\rm C"